# Benchmarking Chinese Commonsense Reasoning of LLMs: From Chinese-Specifics to Reasoning-Memorization Correlations

### v1

@@ Abstract
We introduce CHARM, the first benchmark for comprehensively and in-depth evaluating the commonsense reasoning ability of large language models (LLMs) in Chinese, which covers both globally known and Chinese-specific commonsense. We evaluated 7 English and 12 Chinese-oriented LLMs on CHARM, employing 5 representative prompt strategies for improving LLMs’ reasoning ability, such as Chain-of-Thought. Our findings indicate that the LLM’s language orientation and the task’s domain influence the effectiveness of the prompt strategy, which enriches previous research findings. We built closely-interconnected reasoning and memorization tasks, and found that some LLMs struggle with memorizing Chinese commonsense, affecting their reasoning ability, while others show differences in reasoning despite similar memorization performance. We also evaluated the LLMs’ memorization-independent reasoning abilities and analyzed the typical errors. Our study precisely identified the LLMs’ strengths and weaknesses, providing the clear direction for optimization. It can also serve as a reference for studies in other fields. We will release CHARM at https://github.com/opendatalab/CHARM.
@@ 1 Introduction
Prompt strategies like Chain of Thought (CoT) Wei et al. (2022) can significantly improve LLMs’ reasoning performance Wang et al. (2022, 2023b). Particularly, as the training corpus of LLMs is primarily in English Touvron et al. (2023a), studies Shi et al. (2022); Huang et al. (2023a); Zhang et al. (2023a) have shown that for non-English reasoning tasks, some LLMs perform better when reasoning in English than the native language. We evaluate 7 English and 12 Chinese-oriented LLMs on CHARM, employing 5 representative prompt strategies. The result shows that prompt strategies’ effectiveness depends on the LLMs’ orientation and the benchmark task’s domain, which enriches prior research and guides performance assessment and strategy choice for non-English LLMs.
LLMs’ commonsense reasoning relies on memorization. Exploring the correlation between memorization and reasoning offers insights into LLMs, aiding deeper understanding and suggesting ways to enhance these abilitiesBian et al. (2023). Some benchmarks Yu et al. (2023); Wang et al. (2023a); Fei et al. (2023) aid the research of memorization-reasoning relationships by incorporate tasks for assessing knowledge memorization and application (like reasoning). However, they often use the existing and disparate datasets for different tasks, resulting in a lack of intrinsic connections between these tasks. For instance, the question $Q_{rea}$ tests the LLM’s reasoning with the knowledge piece $K$. However, in memorization tasks, there probably is not any matching questions to determine if the LLM has effectively memorized $K$. Hence, if the LLM fails on $Q_{rea}$, it’s unclear whether due to poor reasoning or forgetfulness of $K$. This results in the disjointed evaluation of memorization and reasoning, failing to uncover their intrinsic links. To address this limitation, we selected suitable reasoning tasks from CHARM’s Chinese domain, and built related memorization questions for each reasoning question (see Figure 1). This design produces the closely-interconnected reasoning and memorization tasks, therefore allows for not only the concurrent evaluation of the two abilities, but also the assessment of memorization-independent reasoning, providing the clear guidance for the LLMs’ enhancement.
@@ 2 Related Work
Commonsense Reasoning Benchmarks There are lots of English commonsense reasoning benchmarks or datasets Davis (2023), some of which have been translated to various languages, including Chinese Conneau et al. (2018); Ponti et al. (2020); Lin et al. (2022). There are also some native Chinese benchmarks related to commonsense reasoning. LogiQA Liu et al. (2020, 2023), sourced from the Chinese civil servant exam questions, evaluates the LLMs’ reading comprehension and logical reasoning across various knowledge types, including basic commonsense knowledges. CLUE Xu et al. (2020), the comprehensive Chinese language understanding benchmark, includes the Natural Language Inference (NLI) tasks, which require the LLMs’ commonsense reasoning abilities. CMMLU Li et al. (2023) is the comprehensive Chinese benchmark that encompasses exam questions from the 67 subjects. It includes tasks that test elementary commonsense and logical reasoning abilities. CORECODE Shi et al. (2023) is the dataset designed for commonsense reasoning and conflict detection in Chinese, presented in the dialogue format.
Prompt Strategy Prompt strategies such as CoT Wei et al. (2022) can effectively boost the reasoning capabilities of LLMs Wang et al. (2022, 2023b). Notably, as the LLM training corpus is primarily in English Touvron et al. (2023a), research reveals that for reasoning tasks in non-English languages, some LLMs exhibit superior performance when reasoning in English as opposed to the native language Shi et al. (2022); Zhang et al. (2023a); Huang et al. (2023a). Kim et al. (2023) proposed a novel cross-language transfer prompt method, which uses both the source and target languages to construct examples.
@@ 3.1 Commonsense Domain
For the two domains, especially for the above 7 aspects, we collected corresponding entities, forming the lists as shown in the Figure 1 and 5. Most of the entities are selected from Gaokao Bench [^1] Gaokao Bench is the collection of China’s university entrance exam questions, which contributes to all the 7 aspects.Zhang et al. (2023b), Douban [^2] https://www.douban.com/ is the popular user-centric cultural review platform in China, which mainly contributes to the Entertainment aspect., Hupu [^3] https://www.hupu.com/ is the large sports community popular in China, which mainly contributes to the Public Figures aspect.. Some entities are collected with the help of searching engines. We only collect the entities that are well-known in China. These entities are then used to create the commonsense reasoning questions, which belong to the corresponding domain and aspect.
@@ 3.2 Reasoning Tasks
When designing the reasoning tasks in CHARM, we bear two criteria in mind. First, the tasks should span both commonsense domains, particularly the 7 Chinese aspects. Second, the global and Chinese tasks should have identical types and settings, differing only in their commonsense domains. From the existing English commonsense reasoning datasets Davis (2023); Suzgun et al. (2022), we selected the following 7 tasks:
@@
The chosen tasks adequately cover both the commonsense domains, particularly the 7 aspects of the Chinese commonsense domain. This coverage enables a comprehensive assessment of LLMs’ commonsense reasoning ability in Chinese. Moreover, the Chinese-domain questions can be created following the similar types and settings as their global counterpart, facilitating the cleaner comparison of the LLMs’ performance across the domains.
All questions in the CHARM reasoning tasks are multiple-choice questions. Detailed information is in Table 1. Question examples of the tasks are in Figure 6 in Appendix B. We use regular expressions to extract the preferred choice from the generation of the LLMs Huang et al. (2023b); Li et al. (2023) and use accuracy as the metric.
3.3 Memorization Task
Shared commonsense knowledge pieces serve as a link between reasoning and memorization questions. From the 7 reasoning tasks, we chose 4 that can be readily associated in this manner, AJ, TU, MMR, SpU, refered as the Memorization-Reasoning-Interconnected (MRI) tasks, and built the reltated memorization questions.
Construction We first extract the commonsense knowledge pieces related to the entities in the corresponding reasoning questions. Information about each entity was collected to the degree sufficient to address the associated reasoning question, and then used to formulate the memorization questions. Following the Knowledge Memorization task in KoLA Yu et al. (2023), we choose free-form QA instead of multiple-choice or true/false questions, which can effectively avoid the impact of randomness. All memorization questions are handcrafted. Question examples are in Figure 7 in Appendix C. The averaged number of related memorization questions for each reasoning question are shown in Table 2.
@@ 3.3 Memorization Task
Judgement and Metric For the MMR task, we use a rule-based matching method for evaluation; for the other three tasks, we use Chat-GPT for judgement. We use accuracy as the metric.
@@
3.4 Construction and Quality Assurance
The question construction of CHARM involves the following three methods:
Handcraft [H] The questions are created mannualy by the authors, based on the entities in §3.1. For reasoning questions in Chinese commonsense domain, we usually refer to the global counter-part and construct them with the same design idea and style, as shown in Figure 6.
Translate [T] We translate the existing English questions and only keep those consistent with the global commonsense domain.
Select [S] We select commonsense reasoning questions from existing Chinese datasets.
Detailed information of the construction of all tasks are shown in Table 1. The distribution of the three methods is shown in Figure 2.
We ensure the quality through these steps: First, two annotators independently verify each question, whether handcrafted or translated. Second, annotators are given only questions, not our answers, and use external resources to respond. They can also flag problematic questions. Last, a question passes if both annotators’ answers match ours; if not, it’s reviewed and modified by multiple authors.
@@ 4.1 Language Models
We evaluated the currently commonly used LLMs, which can be divided into two categories: (1) 7 English LLMs, including GPT series Achiam et al. (2023), LLaMA-2 Touvron et al. (2023b), and Vicuna [^8] https://huggingface.co/lmsys/vicuna-7b-v1.5-16k and https://huggingface.co/lmsys/vicuna-13b-v1.5-16k. (2) 12 Chinese-oriented LLMs, including ChatGLM3999https://huggingface.co/THUDM/chatglm3-6b-32k, Baichuan2 Yang et al. (2023), InternLM2 Team (2023), Yi [^10] https://github.com/01-ai/Yi, DeepSeekBi et al. (2024) and Qwen Bai et al. (2023). For open-source models, we choose the chat version instead of the base version. For closed-source models, we use the official API [^11] We used the gpt-3.5-turbo-1106 version for GPT-3.5 and the gpt-4-1106-preview version for GPT-4.. Detailed information is in Table 3.
@@
AJ
TU
SqU
MMR
SpU
NLI
RC
Avg.
AJ
TU
SqU
MMR
SpU
NLI
RC
Avg.
GPT-3.5-1106
85.33
39.0
65.0
42.0
80.5
61.0
50.5
60.48
90.00
94.0
87.0
46.0
88.5
66.0
49.5
74.43
GPT-4-1106
96.67
60.0
85.0
74.0
86.0
77.0
62.5
77.31 95.33 98.0 97.0 66.0
90.0
72.0
72.0 84.33
LLaMA-2-7B
51.33
36.0
11.0
14.0
49.5
52.0
8.0
31.69
62.67
17.0
14.0
16.0
49.5
22.0
13.0
27.74
LLaMA-2-13B
56.00
33.0
38.0
30.0
58.0
47.0
38.0
42.86
66.67
24.0
39.0
50.0
53.5
57.0
33.5
46.24
LLaMA-2-70B
57.33
37.0
52.0
32.0
55.0
56.0
41.5
47.26
72.67
84.0
73.0
42.0
64.0
61.0
41.5
62.60
Vicuna-7B-v1.5
52.00
29.0
34.0
32.0
51.0
49.0
35.5
40.36
45.33
64.0
37.0
26.0
58.5
52.0
32.5
45.05
Vicuna-13B-v1.5
64.67
25.0
32.0
26.0
51.5
60.0
40.0
42.74
72.67
74.0
41.0
50.0
68.0
61.0
36.0
57.52
ChatGLM3-6B
66.00
40.0
59.0
38.0
77.0
72.0
37.5
55.64
34.00
69.0
71.0
28.0
75.5
63.0
34.0
53.50
Baichuan2-7B
76.00
41.0
48.0
38.0
72.0
53.0
49.5
53.93
55.33
65.0
54.0
26.0
60.5
59.0
29.0
49.83
Baichuan2-13B
85.33
40.0
48.0
46.0
72.5
66.0
51.5
58.48
77.33
74.0
58.0
40.0
71.0
61.0
39.0
60.05
InternLM2-7B
88.00
38.0
58.0
38.0
76.0
81.0
25.0
57.71
74.67
80.0
62.0
20.0
78.0
76.0
23.5
59.17
InternLM2-20B
88.00
55.0
54.0
44.0
74.5
80.0
23.0
59.79
82.67
83.0
61.0
14.0
74.5
72.0
27.0
59.17
Yi-6B
70.67
32.0
47.0
32.0
75.0
50.0
42.0
49.81
79.33
63.0
43.0
14.0
70.5
57.0
33.5
51.48
Yi-34B
96.00
55.0
89.0
76.0
88.5
72.0
51.5
75.43
88.67
92.0
87.0
56.0
89.0
70.0
47.5
75.74
DeepSeek-7B
81.33
34.0
50.0
50.0
79.5
57.0
31.5
54.76
68.00
76.0
47.0
50.0
72.5
59.0
32.5
57.86
DeepSeek-67B
96.67
57.0
83.0
92.0
87.5
77.0
34.5
75.38
90.00
95.0
86.0
22.0
88.0
73.0
39.0
70.43
Qwen-7B
70.67
38.0
55.0
48.0
71.0
57.0
49.5
55.60
74.67
78.0
69.0
50.0
72.5
55.0
36.0
62.17
Qwen-14B
87.33
54.0
77.0
60.0
82.5
66.0
55.0
68.83
84.00
83.0
83.0
44.0
84.5
71.0
40.0
69.93
Qwen-72B
98.00
59.0
91.0
84.0
86.5
84.0 67.5 81.43
94.00
92.0
93.0
64.0
93.0
71.0
63.5
81.50
Table 4: Accuracy on CHARM reasoning tasks. We selected the empirically optimal prompt strategy: XLT for English LLMs and ZH-CoT for Chinese-oriented LLMs. Bold and underline represent the first and second place respectively. Detailed results are in Table 8 and Table 9 of Appendix E.

### v2

@@ Abstract
We introduce CHARM, the first benchmark for comprehensively and in-depth evaluating the commonsense reasoning ability of large language models (LLMs) in Chinese, which covers both globally known and Chinese-specific commonsense. We evaluated 7 English and 12 Chinese-oriented LLMs on CHARM, employing 5 representative prompt strategies for improving LLMs’ reasoning ability, such as Chain-of-Thought. Our findings indicated that the LLM’s language orientation and the task’s domain influence the effectiveness of the prompt strategy, which enriches previous research findings. We built closely-interconnected reasoning and memorization tasks, and found that some LLMs struggle with memorizing Chinese commonsense, affecting their reasoning ability, while others show differences in reasoning despite similar memorization performance. We also evaluated the LLMs’ memorization-independent reasoning abilities and analyzed the typical errors. Our study precisely identified the LLMs’ strengths and weaknesses, providing the clear direction for optimization. It can also serve as a reference for studies in other fields. We will release CHARM at https://github.com/opendatalab/CHARM.
@@
Benchmarks CN-Lang CSR CN-specifics Dual-Domain Rea-Mem
Most benchmarks in Davis (2023)
XNLI, XCOPA, XStoryCloze
LogiQA, CLUE, CMMLU
CORECODE
CHARM (ours)
Table 1: Comparison of commonsense reasoning benchmarks. “CN-Lang” indicates the benchmark is presented in Chinese language. “CSR” means the benchmark is designed to focus on CommonSense Reasoning. “CN-specific” indicates the benchmark includes elements that are unique to Chinese culture, language, regional characteristics, history, etc. “Dual-Domain” indicates the benchmark encompasses both Chinese-specific and global domain tasks, with questions presented in the similar style and format. “Rea-Mem” indicates the benchmark includes closely-interconnected reasoning and memorization tasks.
@@ 1 Introduction
Prompt strategies like Chain of Thought (CoT) Wei et al. (2022) can significantly improve LLMs’ reasoning performance Wang et al. (2022, 2023b). Particularly, as the training corpus of LLMs is primarily in English Touvron et al. (2023a), studies Shi et al. (2022); Huang et al. (2023a); Zhang et al. (2023a) have shown that for non-English reasoning tasks, some LLMs perform better when reasoning in English than the native language. We evaluated 7 English and 12 Chinese-oriented LLMs on CHARM, employing 5 representative prompt strategies. The result showed that prompt strategies’ effectiveness depends on the LLMs’ orientation and the benchmark task’s domain, which enriches prior research and guides performance assessment and strategy choice for non-English LLMs.
LLMs’ commonsense reasoning relies on memorization. Exploring the correlation between memorization and reasoning offers insights into LLMs, aiding deeper understanding and suggesting ways to enhance these abilitiesBian et al. (2023). Some benchmarks Yu et al. (2023); Wang et al. (2023a); Fei et al. (2023) aid the research of memorization-reasoning relationships by incorporate tasks for assessing knowledge memorization and application (like reasoning). However, they used the existing and disparate datasets for different tasks, resulting in a lack of intrinsic connections between these tasks. For instance, the question $Q_{rea}$ tests the LLM’s reasoning with the knowledge piece $K$. However, in memorization tasks, there probably is not any matching questions to determine if the LLM has effectively memorized $K$. Hence, if the LLM fails on $Q_{rea}$, it’s unclear whether due to poor reasoning or forgetfulness of $K$. This results in the disjointed evaluation of memorization and reasoning, failing to uncover their intrinsic links. To address this limitation, we selected suitable reasoning tasks from CHARM’s Chinese domain, and built related memorization questions for each reasoning question (see Figure 1). This design produces the closely-interconnected reasoning and memorization tasks, therefore allows for not only the concurrent evaluation of the two abilities, but also the assessment of memorization-independent reasoning, providing the clear guidance for the LLMs’ enhancement.
@@ 2 Related Work
Commonsense Reasoning Benchmarks There are lots of commonsense reasoning benchmarks, most of them are in English Davis (2023). Some Chinese commonsense reasoning benchmarks are directly translated from English benchmarks Conneau et al. (2018); Ponti et al. (2020); Lin et al. (2022), which lack the Chinese specifics. There are some native Chinese benchmarks that include some Chinese-specific factors and involve commonsense reasoning to a certain extent, such as LogiQA Liu et al. (2020, 2023), CLUE Xu et al. (2020) and CMMLU Li et al. (2023). However, they are not designed for commonsense reasoning, therefore containing the large portion of irrelevant tasks and questions. CORECODE Shi et al. (2023) is the benchmark for Chinese commonsense reasoning and commonsense conflict detection, but it is not strictly designed to distinguish the Chinese-specific and global domains when compared with CHARM. In addition, CHARM has the closely-interconnected reasoning and memorization tasks, which are not included in previous commonsense reasoning benchmarks. The comparison of CHARM with previous commonsense reasoning benchmarks is shown in Table 1.
Prompt Strategy Prompt strategies such as CoT Wei et al. (2022) can effectively boost the reasoning capabilities of LLMs Wang et al. (2022, 2023b). Notably, as the LLM training corpus is primarily in English Touvron et al. (2023a), research revealed that for reasoning tasks in non-English languages, some LLMs exhibit superior performance when reasoning in English as opposed to the native language Shi et al. (2022); Zhang et al. (2023a); Huang et al. (2023a). Kim et al. (2023) proposed a novel cross-language transfer prompt method, which uses both the source and target languages to construct examples.
@@ 3.1 Commonsense Domain
For the two domains, especially for the above 7 aspects, we collected corresponding entities, forming the lists as shown in the Figure 1 and 5. Most of the entities were selected from Gaokao Bench [^1] Gaokao Bench is the collection of China’s university entrance exam questions, which contributes to all the 7 aspects.Zhang et al. (2023b), Douban [^2] https://www.douban.com/ is the popular user-centric cultural review platform in China, which mainly contributes to the Entertainment aspect., Hupu [^3] https://www.hupu.com/ is the large sports community popular in China, which mainly contributes to the Public Figures aspect.. Some entities were collected with the help of searching engines. We only collected the entities that are well-known in China. These entities were then used to create the commonsense reasoning questions, which belong to the corresponding domain and aspect.
@@ 3.2 Reasoning Tasks
When designing the reasoning tasks in CHARM, we beared two criteria in mind. First, the tasks should span both commonsense domains, particularly the 7 Chinese aspects. Second, the global and Chinese tasks should have identical types and settings, differing only in their commonsense domains. From the existing English commonsense reasoning datasets Davis (2023); Suzgun et al. (2022), we selected the following 7 tasks:
@@
The chosen tasks adequately cover both the commonsense domains, particularly the 7 aspects of the Chinese commonsense domain. This coverage enables a comprehensive assessment of LLMs’ commonsense reasoning ability in Chinese. Moreover, the Chinese-domain questions could be created following the similar types and settings as their global counterpart, facilitating the cleaner comparison of the LLMs’ performance across the domains.
All questions in the CHARM reasoning tasks are multiple-choice questions. Detailed information is in Table 2. Question examples of the tasks are in Figure 6 in Appendix B. We used regular expressions to extract the preferred choice from the generation of the LLMs Huang et al. (2023b); Li et al. (2023) and used accuracy as the metric.
3.3 Construction of Reasoning Tasks
The construction of CHARM reasoning tasks involved the following three methods:
Translation [T] was applied to some global domain reasoning tasks. We translated the English commonsense reasoning benchmarks mentioned in §3.2 using GPT-3.5. Then we replaced the English names with commonly used Chinese names and manually screen the translated questions, retaining those without translation errors and accepted as commonsense globally.
Selection [S] We selected the excellent native Chinese datasets, LogiQA Liu et al. (2020), and CLUE Xu et al. (2020), and chose the questions that meet the requirements for CHARM.
Handcraft [H] was mainly applied to the most Chinese domain reasoning tasks. We used the entities in §3.1, and referred to the corresponding global domain task questions (from [T] or [S]) to construct questions with the same type and style. This ensured that the same reasoning task in two domains only differs in the commonsense domain, thus facilitating cleaner comparative analysis, as shown in Figure 6.
Detailed construction information of all reasoning tasks are shown in Table 2.
3.4 Memorization Tasks
Shared commonsense knowledge pieces serve as links between reasoning and memorization questions. From the 7 reasoning tasks, we chose 4 that can be readily associated in this manner, AJ, TU, MMR, SpU, referred as the Memorization-Reasoning-Interconnected (MRI) tasks, and built the related memorization questions.
Construction We first extracted the commonsense knowledge pieces related to the entities in the corresponding reasoning questions. Information about each entity was collected to the degree sufficient to address the associated reasoning question, and then used to formulate the memorization questions. Following the Knowledge Memorization task in KoLA Yu et al. (2023), we chose free-form QA instead of multiple-choice or true/false questions, which can effectively avoid the impact of randomness. All memorization questions were handcrafted [H]. Question examples are in Figure 7 in Appendix C. The averaged number of related memorization questions for each reasoning question are shown in Table 3.
@@ 3.4 Memorization Tasks
Judgement and Metric For the memorization task of MMR, we used a rule-based matching method for evaluation; for the other three tasks, we used GPT-3.5 for judgement. We used accuracy as the metric.
@@
3.5 Quality Assurance
The distribution of the contruction methods for CHARM is shown in Figure 2. After construction of CHARM, we conducted the quality assurance to ensure the quality of the questions. We hired professional NLP annotators to review the questions. The quality assurance process involved five steps: (1) We prepared and assigned annotation task packages to annotators; (2) We trained annotators, emphasizing the avoidance of social bias; (3) We conducted a trial review on a random 20% of questions to fine adjust the review process; (4) Two annotators independently reviewed each question and provided answers without seeing our answers, and a question passed only if both annotators’ answers matched ours and they found no issues with the question; (5) For questions that failed in step 4, authors discussed whether to retain, discard, or correct them based on the nature of the issues identified.
Details about the quality assurance are in Appendix D.
@@ 4.1 Language Models
We evaluated the currently commonly used LLMs, which can be divided into two categories: (1) 7 English LLMs, including GPT series Achiam et al. (2023), LLaMA-2 Touvron et al. (2023b), and Vicuna [^8] https://huggingface.co/lmsys/vicuna-7b-v1.5-16k and https://huggingface.co/lmsys/vicuna-13b-v1.5-16k. (2) 12 Chinese-oriented LLMs, including ChatGLM3999https://huggingface.co/THUDM/chatglm3-6b-32k, Baichuan2 Yang et al. (2023), InternLM2 Team (2023), Yi [^10] https://github.com/01-ai/Yi, DeepSeekBi et al. (2024) and Qwen Bai et al. (2023). For open-source models, we chose the chat version instead of the base version. For closed-source models, we used the official API [^11] We used the gpt-3.5-turbo-1106 version for GPT-3.5 and the gpt-4-1106-preview version for GPT-4.. Detailed information is in Table 4.
We used opencompass [^12] https://github.com/open-compass/opencompass in all our experiments. For all LLMs, the maximum out length was set to 512. For all open-source LLMs, we used the default settings in opencompass: the decoding temperature was the default value of the huggingface transformers library [^13] https://github.com/huggingface/transformers, which is 1.0; do_sample was set to False; the PyTorch numerical type was bf16. For closed-source models (GPT-3.5 and GPT-4), we used the default settings in opencompass: the temperature was set to 0.7.
@@
AJ TU SqU MMR SpU NLI RC Avg. AJ TU SqU MMR SpU NLI RC Avg.
Random Baseline 50.0 25.0 25.0 25.0 50.0 33.3 25.0 33.33 50.0 19.9 25.0 25.0 50.0 33.3 25.0 32.60
GPT-3.5-1106 85.33 39.0 65.0 42.0 80.5 61.0 50.5 60.48 90.00 94.0 87.0 46.0 88.5 66.0 49.5 74.43
GPT-4-1106 96.67 60.0 85.0 74.0 86.0 77.0 62.5 77.31 95.33 98.0 97.0 66.0 90.0 72.0 72.0 84.33
LLaMA-2-7B 51.33 36.0 11.0 14.0 49.5 52.0 8.0 31.69 62.67 17.0 14.0 16.0 49.5 22.0 13.0 27.74
LLaMA-2-13B 56.00 33.0 38.0 30.0 58.0 47.0 38.0 42.86 66.67 24.0 39.0 50.0 53.5 57.0 33.5 46.24
LLaMA-2-70B 57.33 37.0 52.0 32.0 55.0 56.0 41.5 47.26 72.67 84.0 73.0 42.0 64.0 61.0 41.5 62.60
Vicuna-7B-v1.5 52.00 29.0 34.0 32.0 51.0 49.0 35.5 40.36 45.33 64.0 37.0 26.0 58.5 52.0 32.5 45.05
Vicuna-13B-v1.5 64.67 25.0 32.0 26.0 51.5 60.0 40.0 42.74 72.67 74.0 41.0 50.0 68.0 61.0 36.0 57.52
ChatGLM3-6B 66.00 40.0 59.0 38.0 77.0 72.0 37.5 55.64 34.00 69.0 71.0 28.0 75.5 63.0 34.0 53.50
Baichuan2-7B 76.00 41.0 48.0 38.0 72.0 53.0 49.5 53.93 55.33 65.0 54.0 26.0 60.5 59.0 29.0 49.83
Baichuan2-13B 85.33 40.0 48.0 46.0 72.5 66.0 51.5 58.48 77.33 74.0 58.0 40.0 71.0 61.0 39.0 60.05
InternLM2-7B 88.00 38.0 58.0 38.0 76.0 81.0 25.0 57.71 74.67 80.0 62.0 20.0 78.0 76.0 23.5 59.17
InternLM2-20B 88.00 55.0 54.0 44.0 74.5 80.0 23.0 59.79 82.67 83.0 61.0 14.0 74.5 72.0 27.0 59.17
Yi-6B 70.67 32.0 47.0 32.0 75.0 50.0 42.0 49.81 79.33 63.0 43.0 14.0 70.5 57.0 33.5 51.48
Yi-34B 96.00 55.0 89.0 76.0 88.5 72.0 51.5 75.43 88.67 92.0 87.0 56.0 89.0 70.0 47.5 75.74
DeepSeek-7B 81.33 34.0 50.0 50.0 79.5 57.0 31.5 54.76 68.00 76.0 47.0 50.0 72.5 59.0 32.5 57.86
DeepSeek-67B 96.67 57.0 83.0 92.0 87.5 77.0 34.5 75.38 90.00 95.0 86.0 22.0 88.0 73.0 39.0 70.43
Qwen-7B 70.67 38.0 55.0 48.0 71.0 57.0 49.5 55.60 74.67 78.0 69.0 50.0 72.5 55.0 36.0 62.17
Qwen-14B 87.33 54.0 77.0 60.0 82.5 66.0 55.0 68.83 84.00 83.0 83.0 44.0 84.5 71.0 40.0 69.93
Qwen-72B 98.00 59.0 91.0 84.0 86.5 84.0 67.5 81.43 94.00 92.0 93.0 64.0 93.0 71.0 63.5 81.50
Table 5: Accuracy on CHARM reasoning tasks. We selected the empirically optimal prompt strategy: XLT for English LLMs and ZH-CoT for Chinese-oriented LLMs. Bold and underline represent the first and second place respectively. Detailed results are in Table 9 and Table 10 of Appendix F.